Find the general integral of the equation

(x-y)p+(y-x-z)q=z,

And the particular solution through the circle z=1,x2+y2=1


1
Expert's answer
2021-03-31T16:46:50-0400

Comparing with "Pp+Qq=R," we get

"P=x-y, Q=y-x-z, R=z"

The corresponding system of equations is


"\dfrac{dx}{x-y}=\dfrac{dy}{y-x-z}=\dfrac{dz}{z}"

"\dfrac{dx+dy+dz}{x-y+y-x-z+z}"

"\dfrac{dx+dy+dz}{0}=\dfrac{d(x+y+z)}{0}"

"d(x+y+z)=0"

"x+y+z=c_1" is the first independent integral.

Consider


"\dfrac{dx-dy+dz}{x-y-y+x+z+z}"

"\dfrac{dx-dy+dz}{2(x-y+z)}=\dfrac{d(x-y+z)}{2(x-y+z)}"

"\dfrac{d(x-y+z)}{2(x-y+z)}=\dfrac{dz}{z}"

"\dfrac{d(x-y+z)}{x-y+z}=2\cdot\dfrac{dz}{z}"

"\ln(x-y+z)=\ln(z^2)+\ln c_2"

"\dfrac{x-y+z}{z^2}=c_2" is the second independent integral.

The general solution is


"\phi(x+y+z, \dfrac{x-y+z}{z^2} )=0"

Let "x+y+z=t" and "\dfrac{x-y+z}{z^2} =u."


"\phi(t, u)=0"

For "z=1" we get "t=x+y+1" and "u=\dfrac{x-y+1}{1^2}"

Then


"x=\dfrac{t+u}{2}-1, y=\dfrac{t-u}{2}"

Given "x^2+y^2=1"

"(\dfrac{t+u}{2})^2-(t+u)+1+(\dfrac{t-u}{2})^2=1"

"t(t-2)+u(u-2)=0"

Hence


"(x+y+z)(x+y+z-2)"

"+(\dfrac{x-y+z}{z^2})(\dfrac{x-y+z}{z^2}-2)=0"